# You Cannot Feed Two Birds with One Score: the Accuracy-Naturalness Tradeoff in Translation

### v1

@@ Appendix A Selecting distributional divergences
This viewpoint reveals the main practical issue with IPMs, which was already noted for $f$-divergences by Theis (2024): the optimal critic depends on the distributions $Q$ and $P$ that we are trying to evaluate. One solution that is often employed in lossy data compression is therefore to train the critic alongside the the generative model, leading to an adversarial training framework (Mentzer et al., 2020; Theis, 2024). The advantage of this approach is that besides obtaining a critic for evaluating the model at test time, it allows us to optimise our models using the accuracy-naturalness objective in Equation 5 directly. Unfortunately, due to the discrete nature of text outputs, adversarial approaches of text generation have lagged behind autoregressive approaches (Caccia et al., 2020), hence more research is needed before translation models can be trained using the analogous accuracy-naturalness objective. Hence, we focus on metrics that do not require optimisation: no-reference metrics. Following Section 4, we will be interested in quantities that can be estimated using the reference distribution $R_{\mathbf{y}}$ and the system outputs ${\mathbf{Y}}^{m}$.
@@ A.1 Using LLM model scores
where $\mathbb{H}[Q\|P]$ is the cross-entropy of $Q$ from $P$. The issue here, is that $\mathbb{H}[Q_{\mathbf{y}}^{m}]$ could differ wildly among the translation systems, and hence the $C^{m}$ scores are incomparable! A different perspective on this problem is to note that the minimizer of the cross-entropy is in general not $P_{\mathbf{y}}^{LM}$ (unless $P_{\mathbf{y}}^{LM}$ is deterministic, which is very uninteresting):
@@ A.2 Using normalised LLM model scores
An alternative, more feasible solution is to take inspiration from the outlier detection literature, and estimate the scores using a universal source coding algorithm such as Zip (Serrà et al., 2020).  [^7] Technically, should take the minimum across several compressors for better approximation (Appendix C in Serrà et al. (2020)). Concretely, let $\ell^{zip}({\mathbf{y}})$ denote the length of the Zip compressor’s output in bits upon compressing the string ${\mathbf{y}}$. If ${\mathbf{y}}$ is long enough, $\ell^{zip}({\mathbf{y}})$ should converge to $-\log q^{m}({\mathbf{y}}_{n})$, which warrants the estimate
@@
These metrics admit an alternative interpretation as an approximation to a statistical distance which we describe in Section 4; see Appendix E for technical details.
@@ Appendix B A generalisation of Theorem 1 of Blau & Michaeli
A distortion measure $\Delta:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is distribution preserving at $P_{{\mathbf{x}},{\mathbf{y}}}$ with respect to $P_{\mathbf{y}}$ if for the system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ maximising Equation 1 we have $Q_{\mathbf{y}}^{*}=P_{\mathbf{y}}$.
@@
Let $\Delta$ be a distortion metric (see Section 2). We say that the minimizer of $\Delta$ is non-unique if there exist two distributions $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1},Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{2}$ such that
@@
The proof proceeds analogously to the proof of Theorem 1 in Blau & Michaeli (2018). We now proceed to prove Theorem B.1:
@@
Now, since by Equation 16 the expectation with respect to $P_{\mathbf{x}}$ and the minimization are exchangeable, the minimizer $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ does not depend on $P_{\mathbf{x}}$.
@@
Now, since $\Delta$ is distribution preserving at $\tilde{P}_{{\mathbf{x}},{\mathbf{y}}}$ by Fact 2, we must have $\tilde{Q}_{\mathbf{y}}^{*}=\tilde{P}_{\mathbf{y}}$, where $\tilde{Q}_{\mathbf{y}}^{*}$ is the marginal distribution defined by the minimizer $\tilde{Q}_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ of $\Delta$ with respect to $\tilde{P}_{{\mathbf{x}},{\mathbf{y}}}$. This means that Equations 18 and 19 must be equal, hence we must also have
@@
Follows from Lemma B.1 combined with Lemma 2 from Blau & Michaeli (2018). ∎
To finish the proof, assume by way of contradiction that the $P_{{\mathbf{x}},{\mathbf{y}}}$ defines a non-deterministic relationship from ${\mathbf{x}}\to{\mathbf{y}}$ and that the distortion $\Delta$ is stably distribution preserving at $P_{{\mathbf{x}},{\mathbf{y}}}$. Then, by Lemma B.2, the minimizer of $\Delta$ is $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}=P_{{\mathbf{y}}\mid{\mathbf{x}}}$ and this estimator is unique. However, by Lemma B.3, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ is non-unique, which leads to a contradiction. ∎
Appendix C Proof of 3.1
@@
Now note that since $-D$ is bounded from above, for $N\geq N^{\prime}$ we have $-D(Q_{\mathbf{y}},R_{\mathbf{y}})\geq N\geq N^{\prime}$, hence $C_{N}\subseteq C_{N^{\prime}}$. Therefore, $A(N)$ is the supremum of the same objective but over a bigger constraint set, hence we must have $A(N)\leq A(N^{\prime})$.
@@
In other words, $-\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{0})=A(N_{0})$ and $-\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1})=A(N_{1})$. This allows us to rewrite Equation 21 as
@@
Hence, to show Equation 24, it remains to show that
@@
where the inequality follows since $-D(Q_{\mathbf{y}}^{\lambda},R_{\mathbf{y}})=N_{\lambda}$ by definition, hence $-D(Q_{\mathbf{y}},R_{\mathbf{y}})\geq N_{\lambda}$ is in the admissible set. Finally, note that
@@ Appendix D MQM details
For the en $\to$ de and ja $\to$ zh language pairs, the WMT24 evaluation exactly followed the evaluation scheme proposed in Freitag et al. (2021). The error categories are shown in Table 1 (limited to the actual errors found in the data), together with our classification into accuracy or fluency errors. This classification is straightforward, as most categories are already labelled as such. It is worth noting that we ignored the “Source issue” and “Other” categories, as we found the marked errors to be uninformative in most cases. In principle it would be possible to classify most of the individual errors in this category to either adequacy or fluency, but it would involve a huge amount of additional manual work. The en $\to$ es translation used another error schema. Its classification into adequacy and fluency errors is shown in Table 2.
@@
Accuracy/Addition Accuracy/Creative Reinterpretation Accuracy/Gender Mismatch Accuracy/Mistranslation Accuracy/Omission Accuracy/Source language fragment Non-translation! Fluency/Grammar Fluency/Inconsistency Fluency/Punctuation Fluency/Register Fluency/Spelling Fluency/Text-Breaking Locale convention/Address format Locale convention/Currency format Locale convention/Time format Style/Archaic or obscure word choice Style/Bad sentence structure Style/Unnatural or awkward Terminology/Inappropriate for context Terminology/Inconsistent Other Source issue
@@
Addition Agreement Do not translate Mistranslation MT hallucination Omission Untranslated Wrong named entity Wrong term Capitalization Date-time format Inconsistency Lacks creativity Grammar Measurement format Number format Punctuation Register Spelling Unnatural flow Whitespace Word order Wrong language variety Other Source issue
@@
Appendix E Technical details for the divergence in Section 4.1
In this section, we define a generalised variant of the divergence from Section 4.1. Our starting point is to notice that IPMs could be viewed as a certain infinity-norm, since we taking a supremum as part of their definition. This suggests that we should be able to define analogous $p$-norms for $p\in[1,\infty)$ over appropriate spaces of distributions too.
@@
Then, similarly to $\lVert\cdot\rVert_{\mathcal{F}}$, this induces a seminorm on $\mathbb{P}$:
@@ E.1 Examples of $D_{p}$ when $\mathcal{P}$ is of Gaussian process type
for link function $\lambda$ and Gaussian process $\mathcal{GP}(m,k)$ with mean function $m$ and covariance function $k$ and where $\circ$ indicates function composition. It turns out, that in these cases, $D_{p}$ is intimately related to the maximum mean discrepancy $\mathrm{MMD}_{k}$ (Gretton et al., 2012), where the MMD is with respect to a reproducing kernel Hilbert space with the kernel $k$ taken to be the covariance function of the Gaussian process above. Indeed, we show in Section E.1.1 that in certain cases, $D_{p}$ and $\mathrm{MMD}_{k}$ are even identical up to a multiplicative constant. However, as we are averaging instead of maximising in the case of $D_{p}$, in such cases we might call it mean-mean discrepancy (MeMeD).
@@
[…]
Substituting this back, proves the first part of our claim. Now, the fact Equation 38 is an unbiased estimator follows from the theory of U-statistics, similarly to the argument of Lemma 6 of Gretton et al. (2012). ∎
To demonstrate the power of Theorem E.2, we now consider three concrete cases when $\mathcal{P}$ is of Gaussian process type.
@@ E.1.1 When $\mathcal{P}$ is a Gaussian process
The simplest case of $\mathcal{P}$ being of Gaussian process type is when the link function is simply the identity. Indeed, in this case we can compute the entire family of $D_{p}$ metrics for all $p\in[0,\infty)$ without invoking Theorem E.2:
@@
[…]
@@ E.1.2 When $\mathcal{P}$ is a log-Gaussian process
Let $P$ and $Q$ be probability measures over $\mathcal{X}$. Let $\mathcal{P}$ be a log-Gaussian process with mean function $m$ and covariance function $k$, i.e. for $f\sim\mathcal{P}$ we have $\log f\sim\mathcal{GP}(m,k)$. Then, $D_{2}$ is of GP type and can be computed according to Theorem E.2 by setting:
@@
By applying Theorem E.2 with $\lambda=\log$, we simply need to compute
@@
$D_{2}(Q,P\mid\mathcal{P})=e\cdot\left\lvert\mathbb{E}_{X\sim Q}\left[p_{LM}(X)^{-1/\lvert x\rvert}\right]-\mathbb{E}_{X\sim P}\left[p_{LM}(X)^{-1/\lvert x\rvert}\right]\right\rvert$ (58)
@@ E.1.3 When $\mathcal{P}$ is a Probit process
Let $P$ and $Q$ be probability measures over $\mathcal{X}$. Let $\mathcal{P}$ be a Probit process with mean function $m$ and covariance function $k$, i.e. for $f\sim\mathcal{P}$ we have $\Phi^{-1}f\sim\mathcal{GP}(m,k)$. Then, $D_{2}$ is of GP type and can be computed according to Theorem E.2 by letting
@@
Similarly as in the previous section, we can apply Theorem E.2 with $\lambda=\Phi$. Then, we are interested in computing
@@
[…]
Note, that $[U,V]=\mu+\Sigma^{1/2}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$, from which we get $U=a\epsilon_{1}+\mu_{1}$ and $V=b\epsilon_{1}+c\epsilon_{2}+\mu_{2}$, where $a=\sqrt{\Sigma_{11}},b=\Sigma_{21}/a$ and $c=\sqrt{\Sigma_{22}-b^{2}}$ follow from the Cholesky decomposition of $\Sigma$. Hence, we find

### v3

@@ Appendix A Selecting distributional divergences
This viewpoint reveals the main practical issue with IPMs, which was already noted for $f$-divergences by Theis (2024): the optimal critic depends on the distributions $Q$ and $P$ that we are trying to evaluate. One solution that is often employed in lossy data compression is therefore to train the critic alongside the the generative model, leading to an adversarial training framework (Mentzer et al., 2020; Theis, 2024). The advantage of this approach is that besides obtaining a critic for evaluating the model at test time, it allows us to optimise our models using the accuracy-naturalness objective in Equation˜5 directly. Unfortunately, due to the discrete nature of text outputs, adversarial approaches of text generation have lagged behind autoregressive approaches (Caccia et al., 2020), hence more research is needed before translation models can be trained using the analogous accuracy-naturalness objective. Hence, we focus on metrics that do not require optimisation: no-reference metrics. Following Section˜4, we will be interested in quantities that can be estimated using the reference distribution $R_{\mathbf{y}}$ and the system outputs ${\mathbf{Y}}^{m}$.
@@ A.1 Using LLM model scores
where $\mathbb{H}[Q\|P]$ is the cross-entropy of $Q$ from $P$. The issue here, is that $\mathbb{H}[Q_{\mathbf{y}}^{m}]$ could differ wildly among the translation systems, and hence the $C^{m}$ scores are incomparable! A different perspective on this problem is to note that the minimizer of the cross-entropy is in general not $P_{\mathbf{y}}^{LM}$ (unless $P_{\mathbf{y}}^{LM}$ is deterministic, which is uninteresting):
@@ A.2 Using normalised LLM model scores
An alternative, more feasible solution is to take inspiration from the outlier detection literature, and estimate the scores using a universal source coding algorithm such as Zip (Serrà et al., 2020).  [^7] Technically, should take the minimum across several compressors for better approximation (Appendix C in Serrà et al. (2020)). Concretely, let $\ell^{zip}({\mathbf{y}})$ denote the length of the Zip compressor’s output in bits upon compressing the string $\mathbf{y}$}. If $\mathbf{y}$} is long enough, $\ell^{zip}({\mathbf{y}})$ should converge to $\log q^{m}({\mathbf{y}}_{n})$, which warrants the estimate
@@
These metrics admit an alternative interpretation as an approximation to a statistical distance which we describe in Section˜4; see Appendix˜E for technical details.
@@ Appendix B A generalisation of Theorem 1 of Blau & Michaeli
A distortion measure $\Delta:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is distribution preserving at $P_{{\mathbf{x}},{\mathbf{y}}}$ with respect to $P_{\mathbf{y}}$ if for the system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ maximising Equation˜1 we have $Q_{\mathbf{y}}^{*}=P_{\mathbf{y}}$.
@@
Let $\Delta$ be a distortion metric (see Section˜2). We say that the minimizer of $\Delta$ is non-unique if there exist two distributions $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1},Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{2}$ such that
@@
The proof proceeds analogously to the proof of Theorem 1 in Blau & Michaeli (2018). We now proceed to prove Theorem˜B.1:
@@
Now, since by Equation˜16 the expectation with respect to $P_{\mathbf{x}}$ and the minimization are exchangeable, the minimizer $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ does not depend on $P_{\mathbf{x}}$.
@@
Now, since $\Delta$ is distribution preserving at $\tilde{P}_{{\mathbf{x}},{\mathbf{y}}}$ by Fact 2, we must have $\tilde{Q}_{\mathbf{y}}^{*}=\tilde{P}_{\mathbf{y}}$, where $\tilde{Q}_{\mathbf{y}}^{*}$ is the marginal distribution defined by the minimizer $\tilde{Q}_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ of $\Delta$ with respect to $\tilde{P}_{{\mathbf{x}},{\mathbf{y}}}$. This means that Equations˜18 and 19 must be equal, hence we must also have
@@
Follows from Lemma˜B.1 combined with Lemma 2 from Blau & Michaeli (2018). ∎
To finish the proof, assume by way of contradiction that the $P_{{\mathbf{x}},{\mathbf{y}}}$ defines a non-deterministic relationship from ${\mathbf{x}}\to{\mathbf{y}}$ and that the distortion $\Delta$ is stably distribution preserving at $P_{{\mathbf{x}},{\mathbf{y}}}$. Then, by Lemma˜B.2, the minimizer of $\Delta$ is $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}=P_{{\mathbf{y}}\mid{\mathbf{x}}}$ and this estimator is unique. However, by Lemma˜B.3, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ is non-unique, which leads to a contradiction. ∎
Appendix C Proof of ˜3.1
@@
Now note that since D-D is bounded from above, for $N\geq N^{\prime}$ we have $D(Q_{\mathbf{y}},R_{\mathbf{y}})\geq N\geq N^{\prime}$, hence $C_{N}\subseteq C_{N^{\prime}}$. Therefore, $A(N)$ is the supremum of the same objective but over a bigger constraint set, hence we must have $A(N)\leq A(N^{\prime})$.
@@
In other words, $\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{0})=A(N_{0})$ and $\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{1})=A(N_{1})$. This allows us to rewrite Equation˜21 as
@@
Hence, to show Equation˜24, it remains to show that
@@
where the inequality follows since $D(Q_{\mathbf{y}}^{\lambda},R_{\mathbf{y}})=N_{\lambda}$ by definition, hence $D(Q_{\mathbf{y}},R_{\mathbf{y}})\geq N_{\lambda}$ is in the admissible set. Finally, note that
@@ Appendix D MQM details
For the en $\to$ de and ja $\to$ zh language pairs, the WMT24 evaluation exactly followed the evaluation scheme proposed in Freitag et al. (2021). The error categories are shown in Table˜1 (limited to the actual errors found in the data), together with our classification into accuracy or fluency errors. This classification is straightforward, as most categories are already labelled as such. It is worth noting that we ignored the “Source issue” and “Other” categories, as we found the marked errors to be uninformative in most cases. In principle it would be possible to classify most of the individual errors in this category to either adequacy or fluency, but it would involve a huge amount of additional manual work. The en $\to$ es translation used another error schema. Its classification into adequacy and fluency errors is shown in Table˜2.
@@
Accuracy/Addition
Accuracy/Creative Reinterpretation
Accuracy/Gender Mismatch
Accuracy/Mistranslation
Accuracy/Omission
Accuracy/Source language fragment
Non-translation!
Fluency/Grammar
Fluency/Inconsistency
Fluency/Punctuation
Fluency/Register
Fluency/Spelling
Fluency/Text-Breaking
Locale convention/Address format
Locale convention/Currency format
Locale convention/Time format
Style/Archaic or obscure word choice
Style/Bad sentence structure
Style/Unnatural or awkward
Terminology/Inappropriate for context
Terminology/Inconsistent
Other
Source issue
@@
Addition
Agreement
Do not translate
Mistranslation
MT hallucination
Omission
Untranslated
Wrong named entity
Wrong term
Capitalization
Date-time format
Inconsistency
Lacks creativity
Grammar
Measurement format
Number format
Punctuation
Register
Spelling
Unnatural flow
Whitespace
Word order
Wrong language variety
Other
Source issue
@@
Appendix E Technical details for the divergence in Section˜4.1
In this section, we define a generalised variant of the divergence from Section˜4.1. Our starting point is to notice that IPMs could be viewed as a certain infinity-norm, since we taking a supremum as part of their definition. This suggests that we should be able to define analogous $p$-norms for $p\in[1,\infty)$ over appropriate spaces of distributions too.
@@
Then, similarly to $\lVert\cdot\rVert_{\mathcal{F}}$, this induces a semimetric on $\mathbb{P}$:
@@ E.1 Examples of $D_{p}$ when $\mathcal{P}$ is of Gaussian process type
for link function $\lambda$ and Gaussian process $\mathcal{GP}(m,k)$ with mean function $m$ and covariance function $k$ and where $\circ$ indicates function composition. It turns out, that in these cases, $D_{p}$ is intimately related to the maximum mean discrepancy $\mathrm{MMD}_{k}$ (Gretton et al., 2012), where the MMD is with respect to a reproducing kernel Hilbert space with the kernel $k$ taken to be the covariance function of the Gaussian process above. Indeed, we show in Section˜E.1.1 that in certain cases, $D_{p}$ and $\mathrm{MMD}_{k}$ are even identical up to a multiplicative constant. However, as we are averaging instead of maximising in the case of $D_{p}$, in such cases we might call it mean-mean discrepancy (MeMeD).
@@
$(\mathbb{E}_{f\sim\mathcal{P}}[T_{f}]\pi)(x)$ $=\int\mathbb{E}_{f\sim\mathcal{P}}[f(x)f(y)]\,d\pi(y)$ (43)
$=\int C(x,y)\,d\pi(y).$ (44)
Substituting this back, proves the first part of our claim. Now, the fact Equation˜38 is an unbiased estimator follows from the theory of U-statistics, similarly to the argument of Lemma 6 of Gretton et al. (2012). ∎
To demonstrate the power of Theorem˜E.2, we now consider three concrete cases when $\mathcal{P}$ is of Gaussian process type.
@@ E.1.1 When $\mathcal{P}$ is a Gaussian process
The simplest case of $\mathcal{P}$ being of Gaussian process type is when the link function is simply the identity. Indeed, in this case we can compute the entire family of $D_{p}$ metrics for all $p\in[0,\infty)$ without invoking Theorem˜E.2:
@@
Let $P$ and $Q$ be probability measures over $\mathcal{X}$. Let $\mathcal{P}=\mathcal{GP}(m,k)$ be a Gaussian process with mean function $m$ and kernel $k$ over the sample space. Let $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\,dt$ denote the Euler gamma function and $M(a,b,z)$ denote Kummer’s confluent hypergeometric function. Then,
@@ E.1.2 When $\mathcal{P}$ is a log-Gaussian process
Let $P$ and $Q$ be probability measures over $\mathcal{X}$. Let $\mathcal{P}$ be a log-Gaussian process with mean function $m$ and covariance function $k$, i.e. for $f\sim\mathcal{P}$ we have $\log f\sim\mathcal{GP}(m,k)$. Then, $D_{2}$ can be computed according to Theorem˜E.2 by setting:
@@
By applying Theorem˜E.2 with $\lambda=\log$, we simply need to compute
@@
[…]
@@ E.1.3 When $\mathcal{P}$ is a Probit process
Let $P$ and $Q$ be probability measures over $\mathcal{X}$. Let $\mathcal{P}$ be a Probit process with mean function $m$ and covariance function $k$, i.e. for $f\sim\mathcal{P}$ we have $\Phi^{-1}f\sim\mathcal{GP}(m,k)$. Then, $D_{2}$ can be computed according to Theorem˜E.2 by letting
@@
Similarly as in the previous section, we can apply Theorem˜E.2 with $\lambda=\Phi$. Then, we are interested in computing
@@
$\text{where }\mu=[m(x),m(y)]$ $\text{ and }\Sigma=\begin{bmatrix}k(x,x)&k(x,y)\\ k(x,y)&k(y,y)\end{bmatrix}$ (63)
Note, that $U,V]=\mu+\Sigma^{1/2}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$, from which we get $U=a\epsilon_{1}+\mu_{1}$ and $V=b\epsilon_{1}+c\epsilon_{2}+\mu_{2}$, where $a=\sqrt{\Sigma_{11}},b=\Sigma_{21}/a$ and $c=\sqrt{\Sigma_{22}-b^{2}}$ follow from the Cholesky decomposition of $\Sigma$. Hence, we find